OpenStudy (anonymous):

what is 2 whole and 6/27 simplified

OpenStudy (anonymous):

 simplified

OpenStudy (anonymous):

\[2\frac{ ?6 }{ ?72 }\]

OpenStudy (anonymous):

\[2+\frac{ 6 }{ 72 } = \frac{ 144 }{ 72 }+\frac{ 6 }{ 72 }=\frac{ 150 }{ 72 }\]

OpenStudy (anonymous):

no i mean 2 (whole number) and 6 / 72 simplified

OpenStudy (anonymous):

When I read 2 (whole number) and 6/72 simplified that is what I write. And = +

OpenStudy (anonymous):

.2222(recuring) as a decimal anyway 2/9 as a fraction

OpenStudy (anonymous):

that wasnt there

OpenStudy (anonymous):

you talking to me

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

wat wasnt there

OpenStudy (anonymous):

the answer that you gave me wasn't there

OpenStudy (anonymous):

so wat are the anser choices

OpenStudy (anonymous):

a = 1 and 61/99 b = 99/160 c= 3 and 1/8 d = 18/55

OpenStudy (anonymous):

nevermind i dont need it any more i failed 40%

OpenStudy (anonymous):

thanks though

OpenStudy (anonymous):

awww maannn yhu home schooled

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

Wow. I'm not even sure if that is English Tacamry. Aslynn, I'm going to assume that you do need help because it doesn't sound like you co do it. Can you take a picture of the problem and post it on here?

OpenStudy (anonymous):

why yhu worried about how i speak?

OpenStudy (anonymous):

Because it isn't words and that doesn't help. Nor does your accusation and negative implications of home schooling.

OpenStudy (anonymous):

I'm done with my quiz already

OpenStudy (anonymous):

ahahahahaha thats awesome i agree

OpenStudy (anonymous):

Boy shut the flutter up......is that good engglish now

OpenStudy (anonymous):

lol right....stuff like this anoyys me

OpenStudy (anonymous):

me too